If FEED is coded as 88.
What is the code for BREAD WITH BUTTER

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Aditya
 
FEED = (26 – 6 + 1) + (26 – 5 + 1) + (26 – 5 + 1) + (26 – 4 + 1) = 88
 
hence
 
B = (26 – 2+1) = 25
R = (26 – 18 + 1) = 9
E = (26 – 5 + 1) = 22
A = (26  – 1 + 1) = 26
D = (26 – 4 + 1) = 23
 
W = (26 – 23 + 1) = 4
I = (26 – 9 + 1) = 18
T = (2 6 – 20 +1) = 7
H = (26 – 8 + 1) = 19
 
U = (26 – 21 + 1) = 6
 
Now BREAD WITH BUTTER , you can find